### v1

# Using Confidence Scores to Improve Eyes-free Detection of Speech Recognition Errors

Abstract.
Conversational systems rely heavily on speech recognition to interpret and respond to user commands and queries. Nevertheless, recognition errors may occur, which can significantly affect the performance of such systems. While visual feedback can help detect errors, it may not always be practical, especially for people who are blind or low-vision. In this study, we investigate ways to improve error detection by manipulating the audio output of the transcribed text based on the recognizer’s confidence level in its result. Our findings show that selectively slowing down the audio when the recognizer exhibited uncertainty led to a relative increase of 12% in participants’ error detection ability compared to uniformly slowing down the audio.
1. Introduction
In recent years, there have been notable advancements in Automatic Speech Recognition (ASR) technology, enabling eyes-free interaction (Vtyurina et al., 2019; Huang et al., 2022; Ghosh et al., 2020; Fan et al., 2021) and improving accessibility for devices without a visual display (e.g. Amazon Echo, Google Home). Speech recognition can help make interfaces accessible for individuals with motor impairments (Pradhan et al., 2018; Metatla et al., 2019; Bilmes et al., 2005; Nowrin et al., 2022; Wagner et al., 2012) as well as those who are blind (Azenkot and Lee, 2013). While deep learning models have advanced ASR accuracy (Baevski et al., 2020; Liu et al., 2023), real-world ASR performance is often negatively impacted by background noise, speaker variations, and speaker disfluencies (Goldwater et al., 2010). Despite efforts to improve recognition accuracy in noisy environments using large language models (Wang et al., 2022; Weninger et al., 2015), only a modest relative improvement of 5.7% was obtained (Wang et al., 2022). Azenkot and Lee (Azenkot and Lee, 2013) observed that blind users spent a significant amount of time correcting errors when performing speech dictation tasks. Speech error correction involves a two-step process: 1) detecting errors, and 2) correcting errors (Vertanen and Kristensson, 2009). To fully realize the potential of speech recognition technology, accurate error detection and correction are crucial. In this paper, we focus on the first step, error detection.
Prior work has investigated helping users detect and correct errors by providing visual feedback (Fujiwara, 2016; Price and Sears, 2005). However, visual feedback may not be possible for individuals with visual impairments or for sighted users in situations in which they cannot visually attend to their device. Identifying errors in conversational systems without visual feedback can be challenging for a variety of reasons. Firstly, text-to-speech (TTS) audio can be hard to understand, especially when errors involve short or similar-sounding words (Burke et al., 2006). In a study with sighted users (Hong and Findlater, 2018), participants missed approximately 50% of recognition errors when the TTS audio was played at a rate of 200 words per minute (wpm). This difficulty in understanding TTS can be made even worse when the user is in a noisy environment. Finally, errors may occur infrequently, lulling users into trusting the recognizer.
In this study, we examine how users can detect speech recognition errors through audio-only feedback. Similar to the study by Hong and Findlater (Hong and Findlater, 2018), we investigate the impact of various TTS manipulations on users’ ability to detect ASR errors. Hong and Findlater found that the ability to identify errors improved when audio feedback was delivered at a speech rate of 200 wpm, or even slower at 100 wpm, compared to a higher speech rate of 300 wpm. In this work, we investigate if users’ ability to detect errors can be improved by adjusting the audio feedback using the speech recognizer’s confidence score. The confidence score indicates how certain the ASR system is about the accuracy of its result (Gillick et al., 1997). Furthermore, we investigate participants’ ability to detect errors in both common phrases where all words were in-vocabulary, and challenging phrases where at least one word was out-of-vocabulary (e.g. acronyms, proper names).
We studied the effect of confidence scores on eyes-free error detection by testing four audio annotations: default speech rate, slow speech rate, slow speech rate only for low confidence recognitions, and the inclusion of a beep tone for low confidence recognitions. Results showed that slowing down the TTS audio based on confidence score led to 85.3% accurate error detection, outperforming uniform slowing of the audio. Despite the increased audio length resulting from selective slowing based on the confidence score, participants only experienced a slight 7% increase in the time it took them to review errors compared to the default speech rate condition.
2. User Study
The goal of the user study was to investigate if modulating the audio presentation of the speech recognition results based on a recognizer’s confidence score could improve users’ ability to detect errors.
2.1. Participants
We recruited 48 participants (15 female, 32 male) aged between 21 to 68 (M=36, SD=11.5) via Amazon’s Mechanical Turk, an online crowdsourcing platform. Participants were compensated at a rate of $10 (USD) per hour. Participants completed the experiment in 27 minutes on average. All participants self-reported being native English Speakers. 67% of the participants agreed that they frequently used speech interfaces with 22% agreeing that computers had difficulty understanding their speech. Participants rated statements using a 7-point scale with one denoting strongly disagree and seven denoting strongly agree. See Appendix A for the exact questionnaires we used.
2.2. Study Design
We employed a within-subject experimental design with four counterbalanced conditions:
AllNormal — The recognition result was synthesized into speech and played at 200 wpm. This is similar to the default speaking rate of commercial TTS systems.
AllSlow — The result was played at 70% of the default speaking rate, equivalent to 140 wpm.
UncertainSlow — If the confidence score was below a threshold, the result was played at 140 wpm.
UncertainBeep — If the confidence score was below a threshold, a beep tone was played at the beginning followed by the result played at the default speaking rate of 200 wpm.
In the UncertainSlow and UncertainBeep conditions, we slowed down the TTS audio or added a beep tone when the confidence score fell below a threshold. To establish this threshold, we conducted a pilot study with 12 participants. Using the 480 utterances collected from the pilot study, we recognized the audio using Google’s speech-to-text service111https://cloud.google.com/speech-to-text and tested different thresholds measuring two metrics: 1) the true positive rate (TPR), the proportion of utterances containing one more or recognition errors that were correctly identified as having an error, and 2) the false positive rate (FPR), the proportion of utterances with no errors that were incorrectly identified as having an error. We evaluated the trade-off between the TPR and the FPR at different thresholds with a receiver operating characteristic curve. We identified an optimal threshold of 0.93 for the confidence score, which achieved a sensitivity of 0.85 and a specificity of 0.75, indicating this threshold was able to accurately detect a high percentage of errors while avoiding false positives.
2.3. Procedure
Using a web application, participants first signed a consent form and filled up a demographic questionnaire. At the beginning of the study, participants were provided instructions and completed two practice tasks to familiarize themselves with the task. The audio was played at the default speaking rate in the practice trial. At the start of each condition, we provided participants with a description of how the audio annotation worked for the current condition.
Participants recorded a sentence for each task and the recording was transcribed by Google’s speech-to-text service, then synthesized into speech with Google’s TTS service222https://cloud.google.com/text-to-speech. Speech Synthesis Markup Language (SSML) was used in the TTS request to annotate the audio response (slowing speech rate or adding a beep). Following a delay caused by the speech-to-text and TTS processing (averaging around four seconds), participants were presented with the TTS of the recognition result. Participants were not allowed to play back the TTS audio again. Participants were then asked if the reference sentence matched the audio. If they answered no, indicating a speech recognition error, they were asked to locate the incorrect or missing words, as well as any incorrect additional words that may have appeared between two words in the reference sentence (Figure 1). Participants could only detect errors after the audio finished playing, simulating a real-world scenario where users cannot interrupt the system to correct any errors they detect. At the end of the study, participants completed a final questionnaire about their experience.
We selected phrases from a collection of 407 Twitter phrases (Vertanen et al., 2019). This set included 194 common phrases containing all in-vocabulary words and 213 challenging phrases containing at least one out-of-vocabulary word. We used a vocabulary consisting of 100,000 frequent English words. Challenging phrases included proper nouns and abbreviations that might be difficult for the recognizer to transcribe correctly. We used phrases with 5–10 words. Participants were randomly assigned 40 phrases. Each condition included five common phrases and five challenging phrases that were presented in random order.
[…]
3. Results
In total, we collected 1,920 utterances. Google’s recognizer had a word error rate (WER) of 15% on these utterances. Our analysis includes two key measures: the accuracy of the error detection and the detection time. We conducted a one-way repeated measures ANOVA to compare the four conditions. In cases where the normality assumptions were violated (Shapiro-Wilk test, ¡ 0.05), we employed the non-parametric aligned rank test (ART). We used the Wilcoxon signed-rank test to compare the WER between the common phrases and the challenging phrases.
3.1. Error Detection Accuracy
We calculated how often users correctly determined whether the recognition result contained any errors (i.e. by selecting yes or no after hearing the audio). As shown in Figure 2(a), the proportion of correct error detection was higher in UncertainSlow (85%) compared to AllNormal (80%), AllSlow (76%), and UncertainBeep (79%). A non-parametric ART test revealed a significant difference ( = 4.48, = 0.087, = 0.005). Post-hoc pairwise comparisons with Bonferroni correction found a significant difference between UncertainSlow and All Slow conditions ( = 0.002). In contrast to the previous study by Hong and Findlater (Hong and Findlater, 2018) that reported improved error detection with a slow speech rate, our study did not find a significant difference in error detection between AllSlow and AllNormal conditions ( = 0.94). However, our results suggest that slowing down the audio playback only when necessary might help users to better detect the presence of errors compared to uniformly slowing down the playback.
In our study, participants were presented with both common and challenging phrases. The WER was significantly higher at 17% for challenging phrases compared to 12% for common phrases ( = -0.97, ¡ 0.001). This suggests that our approach of using challenging phrases to elicit more recognition errors was effective. Unfortunately, we lacked sufficient data to reliably analyze the impact of different audio annotations on participants’ ability to identify the specific locations of the errors. This was due to not every participant experiencing a sufficient number of recognition errors in each condition. Across all data, the ratio of correctly locating errors to the actual error locations was 49%, with 2% for insertions, 49% for substitutions, and 62% for deletions. Actual substitution and deletion errors were identified by aligning the reference and recognition transcripts using the Levenshtein distance algorithm (Levenshtein et al., 1966). We determined actual insertion errors by manual review. Across all conditions, the ratio of locating errors was 48% for challenging phrases and 52% for common phrases. This indicates participants missed nearly half of the errors in the transcribed text. Particularly, they struggled with identifying insertions, suggesting that detecting and correcting added words may necessitate greater attention.
Two box plots showing results in the four conditions. The first box plot compares the detection accuracy percentages. The second box plot compared the average detection times in seconds.
3.2. Detection Time
We measured the detection time from the end of the audio playback to the participants’ response of either yes or no. Average detection times were similar: 1.98 seconds in AllNormal, 2.03 seconds in All Slow, 1.86 seconds in Uncertain Slow, and 2.26 seconds in Uncertain Beep condition (Figure 2(b)). These differences were not significant ( = 0.69, = 0.014, = 0.56). However, when considering the total time to review errors by adding the audio playback time, a significant difference was observed ( = 14.2, = 0.23, ¡ 0.001). Participants took less time in reviewing errors in AllNormal (4.52 seconds) condition compared to UncertainSlow (4.85 seconds), AllSlow (5.43), and UncertainBeep (5.42 seconds) conditions. Post-hoc pairwise comparisons revealed significant differences between AllSlow and UncertainSlow ( ¡ 0.001) but not between UncertainSlow and AllNormal conditions ( = 0.16). This indicates that selectively slowing down the audio did not significantly increase the overall review time despite the increased audio length.
4. Discussion
In our study, we used four audio annotations to assess participants’ ability to detect errors in their transcribed speech. Our result suggests that providing users with a measure of confidence in the accuracy of the transcription can be helpful for improving error detection. One limitation of our study is that we only considered sighted native English users. Blind users, for example, may have more experience listening to TTS, which could impact their ability to detect errors. Additionally, non-native users who have accents or different pronunciations may also have different experiences with speech recognition technology. Future research should explore how diverse populations interact with speech recognition technology to detect errors.
In the real world, users may be exposed to various types of noises and distractions that could affect their ability to detect errors. Moreover, users may be engaged in other tasks while using voice assistants (e.g. driving or exercising), which could also affect their ability to detect errors. Future studies could investigate how different contexts and tasks impact users’ ability to detect speech recognition errors.
To create a realistic task, we had participants record themselves speaking sentences and used a speech recognizer to transcribe them. Unlike pre-recorded audio which always contains errors, users in real-world scenarios may not always notice small errors if the recognizer is mostly accurate. However, due to our data lacking sufficient recognition errors for each participant and condition, we were unable to analyze participants’ ability to detect the specific locations of these errors. Nevertheless, the inclusion of challenging phrases was successful in significantly increasing the WER compared to common phrases. We suggest future work consider additional ways to ensure sufficient errors in each condition such as 1) a longer or multi-session study with more utterances per condition, 2) adding noise to participant’s audio to increase errors, or 3) injecting forced errors by occasionally presenting the second best recognition result.
5. Conclusion
In conclusion, our study investigated whether audio annotations based on confidence scores helped participants detect speech recognition errors. We found annotating audio based on confidence score did facilitate better detection of error presence. We believe that our results offer valuable insights that could inform the design of more effective error detection and correction systems, thereby improving the accuracy and usability of conversational systems.
Acknowledgements.
This material is based upon work supported by the NSF under Grant No. IIS-1909248.References
- (1)
- Azenkot and Lee (2013) Shiri Azenkot and Nicole B. Lee. 2013. Exploring the use of speech input by blind people on mobile devices. In Proceedings of the 15th International ACM SIGACCESS Conference on Computers and Accessibility. ACM, Bellevue Washington, 1–8. https://doi.org/10.1145/2513383.2513440
- Baevski et al. (2020) Alexei Baevski, Henry Zhou, Abdelrahman Mohamed, and Michael Auli. 2020. wav2vec 2.0: a framework for self-supervised learning of speech representations. In Proceedings of the 34th International Conference on Neural Information Processing Systems (NIPS’20). Curran Associates Inc., Red Hook, NY, USA, 12449–12460.
- Bilmes et al. (2005) Jeff A. Bilmes, Patricia Dowden, Howard Chizeck, Xiao Li, Jonathan Malkin, Kelley Kilanski, Richard Wright, Katrin Kirchhoff, Amarnag Subramanya, Susumu Harada, and James A. Landay. 2005. The vocal joystick: a voice-based human-computer interface for individuals with motor impairments. In Proceedings of the conference on Human Language Technology and Empirical Methods in Natural Language Processing - HLT ’05. Association for Computational Linguistics, Vancouver, British Columbia, Canada, 995–1002. https://doi.org/10.3115/1220575.1220700
- Burke et al. (2006) Moira Burke, Brian Amento, and Philip Isenhour. 2006. Error Correction of Voicemail Transcripts in SCANMail. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Montréal, Québec, Canada) (CHI ’06). Association for Computing Machinery, New York, NY, USA, 339–348. https://doi.org/10.1145/1124772.1124823
- Fan et al. (2021) Jiayue Fan, Chenning Xu, Chun Yu, and Yuanchun Shi. 2021. Just Speak It: Minimize Cognitive Load for Eyes-Free Text Editing with a Smart Voice Assistant. In The 34th Annual ACM Symposium on User Interface Software and Technology. ACM, Virtual Event USA, 910–921. https://doi.org/10.1145/3472749.3474795
- Fujiwara (2016) Kazuki Fujiwara. 2016. Error Correction of Speech Recognition by Custom Phonetic Alphabet Input for Ultra-Small Devices. In Proceedings of the 2016 CHI Conference Extended Abstracts on Human Factors in Computing Systems (San Jose, California, USA) (CHI EA ’16). Association for Computing Machinery, New York, NY, USA, 104–109. https://doi.org/10.1145/2851581.2890380
- Ghosh et al. (2020) Debjyoti Ghosh, Can Liu, Shengdong Zhao, and Kotaro Hara. 2020. Commanding and Re-Dictation: Developing Eyes-Free Voice-Based Interaction for Editing Dictated Text. ACM Transactions on Computer-Human Interaction 27, 4 (Aug. 2020), 1–31. https://doi.org/10.1145/3390889
- Gillick et al. (1997) L. Gillick, Y. Ito, and J. Young. 1997. A probabilistic approach to confidence estimation and evaluation. In 1997 IEEE International Conference on Acoustics, Speech, and Signal Processing, Vol. 2. 879–882 vol.2. https://doi.org/10.1109/ICASSP.1997.596076 ISSN: 1520-6149.
- Goldwater et al. (2010) Sharon Goldwater, Dan Jurafsky, and Christopher D. Manning. 2010. Which words are hard to recognize? Prosodic, lexical, and disfluency factors that increase speech recognition error rates. Speech Communication 52, 3 (March 2010), 181–200. https://doi.org/10.1016/j.specom.2009.10.001
- Hong and Findlater (2018) Jonggi Hong and Leah Findlater. 2018. Identifying Speech Input Errors Through Audio-Only Interaction. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3173574.3174141
- Huang et al. (2022) Jizhou Huang, Haifeng Wang, Shiqiang Ding, and Shaolei Wang. 2022. DuIVA: An Intelligent Voice Assistant for Hands-free and Eyes-free Voice Interaction with the Baidu Maps App. In Proceedings of the 28th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. ACM, Washington DC USA, 3040–3050. https://doi.org/10.1145/3534678.3539030
- Levenshtein et al. (1966) Vladimir I Levenshtein et al. 1966. Binary codes capable of correcting deletions, insertions, and reversals. In Soviet physics doklady, Vol. 10. Soviet Union, 707–710.
- Liu et al. (2023) Alexander H. Liu, Wei-Ning Hsu, Michael Auli, and Alexei Baevski. 2023. Towards End-to-End Unsupervised Speech Recognition. In 2022 IEEE Spoken Language Technology Workshop (SLT). 221–228. https://doi.org/10.1109/SLT54892.2023.10023187
- Metatla et al. (2019) Oussama Metatla, Alison Oldfield, Taimur Ahmed, Antonis Vafeas, and Sunny Miglani. 2019. Voice User Interfaces in Schools: Co-designing for Inclusion with Visually-Impaired and Sighted Pupils. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–15. https://doi.org/10.1145/3290605.3300608
- Nowrin et al. (2022) Sadia Nowrin, Patricia OrdóñEz, and Keith Vertanen. 2022. Exploring Motor-impaired Programmers’ Use of Speech Recognition. In The 24th International ACM SIGACCESS Conference on Computers and Accessibility. ACM, Athens Greece, 1–4. https://doi.org/10.1145/3517428.3550392
- Pradhan et al. (2018) Alisha Pradhan, Kanika Mehta, and Leah Findlater. 2018. ”Accessibility Came by Accident”: Use of Voice-Controlled Intelligent Personal Assistants by People with Disabilities. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3173574.3174033
- Price and Sears (2005) Kathleen J. Price and Andrew Sears. 2005. Speech-based text entry for mobile handheld devices: An analysis of efficacy and error correction techniques for server-based solutions. International Journal of Human-Computer Interaction 19, 3 (2005), 279–304. https://doi.org/10.1207/s15327590ijhc1903_1 Funding Information: The authors thank Aether Systems, Inc. for their support of this research. This material is based upon work supported by the National Science Foundation under Grants IIS–9910607 and IIS–0121570. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation (NSF)..
- Vertanen et al. (2019) Keith Vertanen, Dylan Gaines, Crystal Fletcher, Alex M. Stanage, Robbie Watling, and Per Ola Kristensson. 2019. VelociWatch: Designing and Evaluating a Virtual Keyboard for the Input of Challenging Text. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–14. https://doi.org/10.1145/3290605.3300821
- Vertanen and Kristensson (2009) Keith Vertanen and Per Ola Kristensson. 2009. Automatic selection of recognition errors by respeaking the intended text. In 2009 IEEE Workshop on Automatic Speech Recognition & Understanding. 130–135. https://doi.org/10.1109/ASRU.2009.5373347
- Vtyurina et al. (2019) Alexandra Vtyurina, Adam Fourney, Meredith Ringel Morris, Leah Findlater, and Ryen W. White. 2019. Bridging Screen Readers and Voice Assistants for Enhanced Eyes-Free Web Search. In The World Wide Web Conference. ACM, San Francisco CA USA, 3590–3594. https://doi.org/10.1145/3308558.3314136
- Wagner et al. (2012) Amber Wagner, Ramaraju Rudraraju, Srinivasa Datla, Avishek Banerjee, Mandar Sudame, and Jeff Gray. 2012. Programming by voice: a hands-free approach for motorically challenged children. In CHI ’12 Extended Abstracts on Human Factors in Computing Systems. ACM, Austin Texas USA, 2087–2092. https://doi.org/10.1145/2212776.2223757
- Wang et al. (2022) Yiming Wang, Jinyu Li, Heming Wang, Yao Qian, Chengyi Wang, and Yu Wu. 2022. Wav2vec-Switch: Contrastive Learning from Original-Noisy Speech Pairs for Robust Speech Recognition. In ICASSP 2022 - 2022 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). 7097–7101. https://doi.org/10.1109/ICASSP43922.2022.9746929 ISSN: 2379-190X.
- Weninger et al. (2015) Felix Weninger, Hakan Erdogan, Shinji Watanabe, Emmanuel Vincent, Jonathan Le Roux, John R. Hershey, and Björn Schuller. 2015. Speech Enhancement with LSTM Recurrent Neural Networks and its Application to Noise-Robust ASR. In Latent Variable Analysis and Signal Separation (Lecture Notes in Computer Science), Emmanuel Vincent, Arie Yeredor, Zbyněk Koldovský, and Petr Tichavský (Eds.). Springer International Publishing, Cham, 91–99. https://doi.org/10.1007/978-3-319-22482-4_11
Appendix A Questionnaire
Figure 3 shows the questions we asked participants at the start of the study. Figure 4 show the questions we asked participants at the end of the study.
[…]